# Random-Set Neural Networks (RS-NN)

### v2

@@ A.2 Statistical guarantees
Firstly, RS-NN can be employed as the ‘underlying model’ in an inductive conformal learning framework (https://cml.rhul.ac.uk/copa2017/presentations/CP_Tutorial_2017.pdf), which builds an empirical cumulative distribution of the ‘non-conformity’ scores of a set of calibration samples, and at test time outputs the set of labels whose empirical CDF is above a desired significance level $\epsilon$ (e.g., 95%).
@@
(i.e., a label $c$ is ‘non-conformal’ if its predicted plausibility, which is defined as $Pl(A)=1-Bel(\Theta\setminus A)$ and has the semantic of an upper probability bound Shafer76 , is low), and compute predictive regions in the usual way:
@@
Experiments on conformal guarantees. We conducted experiments on CIFAR-10 dataset to provide conformal prediction guarantees with non-conformity scores as given in (10). The goal is to assess the reliability of classification predictions by determining a threshold of non-conformity scores that covers 95% of the data.
@@
This prediction threshold is determined by finding the percentile corresponding to $1-\alpha$, where $\alpha$ is set to $0.05$ to achieve 95% coverage. Fig. 8(a) shows the non-conformity scores for all the calibration data, $j=1000$ angelopoulos2021gentle (using the following split: 40000:10000:9000:1000 training, testing, validation and calibration data points respectively for CIFAR-10), with a cut-off threshold that ensures 95% coverage.
@@
Figure 8: (a) Conformal prediction threshold for CIFAR-10 indicating the boundary beyond which predictions are considered non-conformal, (b) Coverage per class for CIFAR-10 dataset, representing the proportion of instances in each class that fall within the predictive regions, (c) Average size of the predictive sets generated for each class of CIFAR-10.
Figs. 8(b) and 8(c) show the coverage and average set size for each class on the test dataset. Coverage is the proportion of prediction sets that actually contain the true outcome, and average set size is the average number of predicted classes per instance. Higher numbers represent more overlap between the classification regions of different classes.
Summarizing the results, (1) the coverage averages across classes over the desired threshold. This indicates that the conformal prediction method is providing reliable prediction intervals that contain the true class label with the specified confidence level; (2) The spread or variability in set sizes across different samples provides insights into how well the prediction sets adapt to the difficulty of the samples. Ideally, the prediction sets should dynamically adjust based on the difficulty of each example, with larger sets for more challenging inputs and smaller sets for easier ones. Fig. 8(c) shows that the average set size for classes 1, 2, 3, 4, and 6 is considerably larger than for the rest, indicating that the model found samples from these as challenging inputs.
@@
Finally, an intriguing alternative approach (and one that is more achievable in the short term) is the study of confidence intervals in a belief functions representation, such as the one we employ in RS-NNs. In fact, recent studies have been looking at extending the notion of confidence interval (https://en.wikipedia.org/wiki/Confidence_interval) to belief functions, under the name of confidence structures (https://hal.science/hal-01576356v3), which generalise standard confidence distributions and generate “frequency-calibrated" belief functions.
Liu and Martin, in particular, have developed an Inferential Model (IM) approach which produces belief functions with well-defined frequentist properties martin2013inferential ; martin2015inferential . An alternative approach relies on the notion of “predictive" belief function denoeux2006constructing , which, under repeated sampling, is less committed than the true probability distribution of interest with some prescribed probability.
A.3 RS-NN Learning Mechanism
Recall from Sec. 3.1 that the ground-truth for a Random Set-CNN is the belief-encoded vector $\mathbf{bel}=\{Bel(A),A\in\mathbb{P}(\mathcal{C})\}$, where $Bel(A)$ is the belief function of a focal set $A$ in the power set $\mathbb{P}$ of classes $\mathcal{C}$. In our method, $Bel(A)$ is 1 if the true class is in subset $A$ and 0 otherwise. Consequently, the belief-encoded ground truth $\mathbf{bel}$ will include multiple occurrences of 1. For instance, in a digit-classification task such as MNIST, if $\{3\}$ is the true class, the belief-encoded ground truth would contain 1s in all sets where $\{3\}$ is present, such as {1, 3}, {0, 3}, {1, 2, 3}, and so forth. This formulation closely resembles multi-label classification.
A.4 Pignistic probability and credal prediction
Figure 9: Upper and lower bounds (9), in dotted red, to the probability of the predicted most likely class according to the pignistic prediction (in solid red) for 100 samples of CIFAR-10.
For additional illustration, the lower and upper bounds (9) to the probability of the top predicted class are plotted in Fig. 9, together with the pignistic probability, for 100 samples of CIFAR-10. The bounds (9) can be efficiently computed from the finite number of vertices of the credal prediction (cuzzolin2008credal , Fig. 4).
@@ Appendix B Related Work
Researchers have recently taken some tentative steps to model uncertainty in machine learning and artificial intelligence. Scientists have adopted measures of epistemic uncertainty kendall2017uncertainties to refuse or delay making a decision geifman2017selective , take actions specifically directed at reducing uncertainty (as in active learning aggarwal2014active ), or exclude highly-uncertain data at decision making time kendall2017uncertainties .
@@
Imprecise probability. Significant work has been done in a credal set setting. Notably, Zaffalon ; Zaffalon2 have proposed the Naive Credal Classifier (NCC) as an extension of the naive Bayes classifier to credal sets, where imprecise probabilities are included in models in the form of sets of classes. ANTONUCCI2017320 have presented graphical models which generalise NCCs to multilabel data. Expected utility maximisation algorithms, such as Bayes-optimal prediction Mortier , and classification with reject option for risk aversion ijcai2018-706 are also based on set-valued predictions. Classification with partial data has been studied by various authors 10.1007/11518655_80 , while a large number of papers have been published on decision trees in the belief function framework elouedi . Significant work in the neural network area was conducted in denoeux and 10.1007/11829898_18 . Classification approaches based on rough sets were proposed in Trabelsi2011ClassificationSB . Ensemble classification burger is another area in which uncertainty theory has been quite impactful, as the problem is one of combining the results of different classifiers. Important contributions have been made by 155943 and ROGOVA1994777 . Regression, on the other hand, has only recently been considered by belief theorists Laanaya2010338 ; 7885577 ; https://doi.org/10.48550/arxiv.2208.00647 . For tasks like image classification in large datasets, most of these approaches have low performance metrics including training time and test accuracy. The epistemic deep learning approach utilizes standard networks to achieve better performance.
@@
Evidential approaches. Within a proper epistemic setting, a significant amount of work has been done by Denoeux and co-authors, and Liu et al. Liu2012291 , on unsupervised learning and clustering in particular in the belief function framework. Quite a lot of work has been done on ensemble classification in the evidential framework 155943 (in particular for neural networks ROGOVA1994777 ), decision trees elouedi00decision , K-nearest neighbour classifiers Denoeux2008classic , and more recently on evidential deep learning classifiers able to quantify uncertainty tong2021evidential . Tong et al.tong2021evidential proposes a convolutional neural network based on Dempster-Shafer theory called the evidential deep-classifier employs utility functions from decision theory to assign utilities on mass functions derived from input features to produce set-valued observations. Another recent approach by Sensoy et al.sensoy proposes an evidential deep learning classifier to estimate uncertainty in the Dirichlet representation. This work is based on subjective logic and learning to form subjective opinions by minimizing the Kullback-Leibler divergence over a uniform Dirichlet distribution. However, epistemic deep learning does not depend on evidence-based parameters for learning.
@@
Conformal prediction. Conformal prediction vovk2005algorithmic provides a framework for estimating uncertainty shafer2008tutorial by applying a threshold on the error the model can make to produce prediction sets, irrespective of the underlying prediction problem. Different variants of conformal predictors are described in papers by Saunders et al. Saunders1999TransductionWC , Nouretdinov et al. Nouretdinov01ridgeregression , Proedrou et al. 10.1007/3-540-36755-1_32 and Papadopoulos et al. 10.5555/1712759.1712773 . Since the computational inefficiency of conformal predictors posed a problem for their use in neural networks, Inductive Conformal Predictors (ICPs) were proposed by Papadopoulos et al. 10.1007/3-540-36755-1_29 ICP . Venn Predictors NIPS2003_10c66082 , cross-conformal predictors crossconformal and Venn-Abers predictors https://doi.org/10.48550/arxiv.1211.0025 were introduced in distribution-free uncertainty quantification using conformal learning.
@@ C.1 Algorithm for Budgeting
Next, we fit a Gaussian Mixture Model (GMM) to the reduced feature vectors of each class. Using the eigenvectors and eigenvalues of the covariance matrix $\Sigma_{c}$ and the mean vector $\mu_{c}$ for each class $c$, we define an ellipsoid covering 95% of the data. The lengths of the ellipsoid’s principal axes are computed as $length_{c,i}=2\sqrt{7.815\lambda_{i}}$, where $\lambda_{i}$ is the $i^{th}$ eigenvalue. The scalar 7.815 corresponds to a 95The class ellipsoids are plotted in a 3D space and the overlap of each subset in the power set of $N$ is computed. As it is computationally infeasible to compute overlap for all $2^{N}$ subsets, we start doing so from cardinality 2 and use early stopping when increasing the cardinality further does not alter the list of most-overlapping sets of classes. We choose the top-$K$ subsets ($A_{K}$) with the highest overlapping ratio, computed as the intersection over union for each subset.
@@ C.2 Algorithm for ECE
0:  confidences: List or array of confidence scores predicted by the model  predictions: List or array of predicted class labels  true_labels: List or array of true class labels  $B$: Number of bins for binning confidence scores
0:  ECE: Expected Calibration Error
1:  Step 1: Normalize Confidences  Normalize the confidence scores to ensure they are in the range [0, 1].
2:  Step 2: Binning  Divide the confidence scores into num_bins bins.
3:  Step 3: Initialize Arrays  Initialize arrays bin_accuracy, bin_confidence, and weights with zeros.
4:  Step 4: Populate Arrays
5:  for each bin do
6:     Identify instances falling into the bin
7:     Calculate mean accuracy and mean confidence within the bin
8:     Update bin_accuracy, bin_confidence, and weights
9:  end for
10:  Step 5: Calculate ECE  Calculate the Expected Calibration Error using the populated arrays.
@@ C.3 Algorithm for AUROC, AUPRC
AUROC (Area Under the Receiver Operating Characteristic Curve) and AUPRC (Area Under the Precision-Recall Curve) are evaluation metrics commonly used to assess the performance of binary classification models, providing insights into their ability to distinguish between positive and negative instances. In the context of evaluating uncertainty estimation in machine learning models, these metrics quantify how well the model separates iD samples from OoD samples. AUROC is derived from the Receiver Operating Characteristic (ROC) curve, which illustrates the trade-off between True Positive Rate (TPR) and False Positive Rate (FPR) across various classification thresholds. The AUROC value represents the area under this curve and ranges from 0 to 1, with higher values indicating better discriminative performance. AUPRC is based on the Precision-Recall curve, which plots Precision against Recall at different classification thresholds. Precision measures the accuracy of positive predictions, while Recall quantifies the ability to capture all positive instances. AUPRC calculates the area under this curve and provides a complementary perspective, particularly valuable when dealing with imbalanced datasets.
Algorithm 3 Algorithm for AUROC, AUPRC
0:  uncertainty_iid: Uncertainty scores for in-distribution samples. uncertainty_ood: Uncertainty scores for out-of-distribution samples.

### v3

@@ A.4 Statistical guarantees
Plugging RS-NN into conformal learning. Indeed, RS-NN can be employed as the ‘underlying model’ in an inductive conformal learning framework [^3] https://cml.rhul.ac.uk/copa2017/presentations/CP_Tutorial_2017.pdf, which builds an empirical cumulative distribution of the ‘non-conformity’ scores of a set of calibration samples, and at test time outputs the set of labels whose empirical CDF is above a desired significance level $\epsilon$ (e.g., 95%).
@@
(i.e., a label $c$ is ‘non-conformal’ if its predicted plausibility, which is defined as $Pl(A)=1-Bel(\Theta\setminus A)$ and has the semantic of an upper probability bound (Shafer, 1976), is low), and compute predictive regions in the usual way:
@@
Experiments on conformal guarantees. Thus, we conducted experiments on CIFAR-10 dataset to provide conformal prediction guarantees with non-conformity scores as given in Eq. 10. The goal is to assess the reliability of classification predictions by determining a threshold of non-conformity scores that covers 95% of the data.
@@
This prediction threshold is determined by finding the percentile corresponding to $1-\alpha$, where $\alpha$ is set to $0.05$ to achieve 95% coverage. Fig. 11 shows the non-conformity scores for all the calibration data, $j=1000$ (Angelopoulos & Bates, 2021) (using the following split: 40000:10000:9000:1000 training, testing, validation and calibration data points respectively for CIFAR-10), with a cut-off threshold that ensures 95% coverage.
@@
Figure 9: Conformal prediction threshold for CIFAR-10 indicating the boundary beyond which predictions are considered non-conformal.
Figure 10: Coverage per class of RS-NN for CIFAR-10 dataset.
Figure 11: Average size of the predictive sets generated for each class of CIFAR-10.
Figs. 11 and 11 show the coverage and average set size for each class on the test dataset. Coverage is the proportion of prediction sets that actually contain the true outcome, and average set size is the average number of predicted classes per instance. Higher numbers represent more overlap between the classification regions of different classes.
Summarizing the results, (1) the coverage averages across classes over the desired threshold. This indicates that the conformal prediction method is providing reliable prediction intervals that contain the true class label with the specified confidence level; (2) The spread or variability in set sizes across different samples provides insights into how well the prediction sets adapt to the difficulty of the samples. Ideally, the prediction sets should dynamically adjust based on the difficulty of each example, with larger sets for more challenging inputs and smaller sets for easier ones. Fig. 11 shows that the average set size for classes 1, 2, 3, 4, and 6 is considerably larger than for the rest, indicating that the model found samples from these as challenging inputs.
@@
Finally, an intriguing alternative approach (and one that is more achievable in the short term) is the study of confidence intervals in a belief functions representation, such as the one we employ in RS-NNs. In fact, recent studies have been looking at extending the notion of confidence interval (https://en.wikipedia.org/wiki/Confidence_interval) to belief functions, under the name of confidence structures (https://hal.science/hal-01576356v3), which generalise standard confidence distributions and generate “frequency-calibrated” belief functions.
Liu and Martin, in particular, have developed an Inferential Model (IM) approach which produces belief functions with well-defined frequentist properties (Martin & Liu, 2013; 2015). An alternative approach relies on the notion of “predictive” belief function (Denœux, 2006), which, under repeated sampling, is less committed than the true probability distribution of interest with some prescribed probability.
@@ Appendix B Related Work
Researchers have recently taken some tentative steps to model uncertainty in machine learning and artificial intelligence. Scientists have adopted measures of epistemic uncertainty (Kendall & Gal, 2017) to refuse or delay making a decision (Geifman & El-Yaniv, 2017), take actions specifically directed at reducing uncertainty (as in active learning (Aggarwal et al., 2014)), or exclude highly-uncertain data at decision making time (Kendall & Gal, 2017).
Imprecise probability. Significant work has been done in a credal set setting. Notably, (Zaffalon, 2002; Corani & Zaffalon, 2008) have proposed the Naive Credal Classifier (NCC) as an extension of the naive Bayes classifier to credal sets, where imprecise probabilities are included in models in the form of sets of classes. (Antonucci & Corani, 2017) have presented graphical models which generalise NCCs to multilabel data. Expected utility maximisation algorithms, such as Bayes-optimal prediction (Mortier et al., 2021), and classification with reject option for risk aversion (Nguyen et al., 2018) are also based on set-valued predictions.
@@
Classification with partial data has been studied by various authors (Vannoorenberghe & Smets, 2005), while a large number of papers have been published on decision trees in the belief function framework (Elouedi1 et al., 2000). Significant work in the neural network area was conducted in (Denoeux, 2000) and (Fay et al., 2006). Classification approaches based on rough sets were proposed in (Trabelsi et al., 2011). Ensemble classification (Burger et al., 2006) is another area in which uncertainty theory has been quite impactful, as the problem is one of combining the results of different classifiers. Important contributions have been made by (Xu et al., 1992) and (Rogova, 2008). Regression, on the other hand, has only recently been considered by belief theorists (Laanaya et al., 2010; Gong & Cuzzolin, 2018; Denœux, 2022). For tasks like image classification in large datasets, most of these approaches have low performance metrics including training time and test accuracy.
@@
Evidential approaches. Within a proper epistemic setting, a significant amount of work has been done by Denoeux and co-authors, and Liu et al. (2012), on unsupervised learning and clustering in particular in the belief function framework. Quite a lot of work has been done on ensemble classification in the evidential framework (Xu et al., 1992) (in particular for neural networks (Rogova, 2008)), decision trees (Elouedi et al., Madrid, 2000), K-nearest neighbour classifiers (Denœux, 2008), and more recently on evidential deep learning classifiers able to quantify uncertainty (Tong et al., 2021). Tong et al.(Tong et al., 2021) proposes a convolutional neural network based on Dempster-Shafer theory called the evidential deep-classifier employs utility functions from decision theory to assign utilities on mass functions derived from input features to produce set-valued observations. Another recent approach by Sensoy et al.(Sensoy et al., 2018) proposes an evidential deep learning classifier to estimate second-order uncertainty in the Dirichlet representation. This work is based on subjective logic and learning to form subjective opinions by minimizing the Kullback-Leibler divergence over a uniform Dirichlet distribution.
These methods represent predictions as Dirichlet distributions, as explored in various works over the last few years (Sensoy et al., 2018; Malinin & Gales, 2018; 2019; Malinin et al., 2019; Charpentier et al., 2020). However, many commonly used loss functions for these networks are flawed, as they fail to ensure that epistemic uncertainty diminishes with more data, violating basic asymptotic assumptions (Bengs et al., 2022). Additionally, some approaches require out-of-distribution (OoD) data for training, which may not always be available and doesn’t guarantee robustness against all types of OoD data. Studies (Ulmer et al., 2023) have shown that OoD detection degrades in certain models under adversarial conditions, and even techniques like normalizing flows (NFs) in posterior networks (Charpentier et al., 2020), while effective, can struggle with OoD data when relying on learned features (Kopetzki et al., 2021; Stadler et al., 2021).
@@
Conformal prediction. Conformal prediction (Vovk et al., 2005) provides a framework for estimating uncertainty (Shafer & Vovk, 2008) by applying a threshold on the error the model can make to produce prediction sets, irrespective of the underlying prediction problem. Different variants of conformal predictors are described in papers by Saunders et al. (Saunders et al., 1999), Nouretdinov et al. (Nouretdinov et al., 2001), Proedrou et al. (Proedrou et al., 2002) and Papadopoulos et al. (Papadopoulos et al., 2008).
Since the computational inefficiency of conformal predictors posed a problem for their use in neural networks, Inductive Conformal Predictors (ICPs) were proposed by Papadopoulos et al. (Papadopoulos et al., 2002a) (Papadopoulos et al., 2002b). Venn Predictors (Vovk et al., 2003), cross-conformal predictors (Vovk, 2012) and Venn-Abers predictors (Vovk & Petej, 2012) were introduced in distribution-free uncertainty quantification using conformal learning. As we show above, RS-NN is compatible with conformal prediction, as a possible underlying model. In the future we plan to extend our random-set representation to conformal learning, by replacing cumulative distribution functions with random sets.
Deterministic uncertainty quantification. Deterministic Uncertainty Quantification (DUQ) methods estimate epistemic uncertainty by analyzing latent representations or using distance-sensitive functions instead of softmax (Alemi et al., 2018; Wu & Goodman, 2020; Liu et al., 2020; Mukhoti et al., 2021; Van Amersfoort et al., 2020). These methods have been applied to areas like object detection (Gasperini et al., 2021) but focus mainly on OoD detection, overlooking calibration — how well uncertainty reflects model performance under shifting distributions. Calibration is crucial for safe deployment but remains underexplored, as DUQs are typically tested only on simple tasks and datasets, leaving their effectiveness in more complex scenarios unproven (Postels et al., 2021).
Uncertainty estimates are poorly calibrated under distributional shifts, especially when compared to scalable Bayesian methods. Regularization techniques based on feature space distances, such as bi-Lipschitz regularization (used by most of these models), do not effectively improve OoD detection or calibration (Postels et al., 2021). This is due to the limitations of distance metrics in high-dimensional data, such as images. Additionally, DUQ methods differ from RS-NN by modeling epistemic uncertainty in the input space rather than the prediction space. They assess uncertainty based on whether the model’s input is an iD or OoD sample. Because of this difference in how epistemic uncertainty is modeled, we cannot directly compare RS-NN to DUQ methods.
@@ Appendix C Algorithms
This section outlines the algorithms integral to the implementation and evaluation of budgeting (§C.1), expected calibration error (ECE) (§C.2), area under the receiver operating characteristic curve (AUROC) and area under the precision-recall curve (AUPRC) (§C.3).
@@ C.1 Algorithm for Budgeting
Next, we fit a Gaussian Mixture Model (GMM) to the reduced feature vectors of each class. Using the eigenvectors and eigenvalues of the covariance matrix $\Sigma_{c}$ and the mean vector $\mu_{c}$ for each class $c$, we define an ellipsoid covering 95% of the data. The lengths of the ellipsoid’s principal axes are computed as $length_{c,i}=2\sqrt{7.815\lambda_{i}}$, where $\lambda_{i}$ is the $i^{th}$ eigenvalue. The scalar 7.815 corresponds to a 95% confidence interval for a chi-square distribution with 3 degrees of freedom. The class ellipsoids are plotted in a 3D space and the overlap of each subset in the power set of $N$ is computed. As it is computationally infeasible to compute overlap for all $2^{N}$ subsets, we start doing so from cardinality 2 and use early stopping when increasing the cardinality further does not alter the list of most-overlapping sets of classes. We choose the top-$K$ subsets ($A_{K}$) with the highest overlapping ratio, computed as the intersection over union for each subset.
@@ C.2 Algorithm for ECE
1:confidences: List or array of confidence scores predicted by the model  predictions: List or array of predicted class labels  true_labels: List or array of true class labels  $B$: Number of bins for binning confidence scores
2:ECE: Expected Calibration Error
3:Step 1: Normalize Confidences  Normalize the confidence scores to ensure they are in the range [0, 1].
4:Step 2: Binning  Divide the confidence scores into num_bins bins.
5:Step 3: Initialize Arrays  Initialize arrays bin_accuracy, bin_confidence, and weights with zeros.
6:Step 4: Populate Arrays
7:for each bin do
8:     Identify instances falling into the bin
9:     Calculate mean accuracy and mean confidence within the bin
10:     Update bin_accuracy, bin_confidence, and weights
11:end for
12:Step 5: Calculate ECE  Calculate the Expected Calibration Error using the populated arrays.
@@ C.3 Algorithm for AUROC, AUPRC
Algorithm 3 Algorithm for AUROC, AUPRC
1:uncertainty_iid: Uncertainty scores for in-distribution samples. uncertainty_ood: Uncertainty scores for out-of-distribution samples.
[…]
4:auroc: Area Under the Receiver Operating Characteristic curve.
5:auprc: Area Under the Precision-Recall curve.
6:Step 1: Concatenate uncertainties;
[…]
8:Step 2: Create and combine labels;
[…]
10:Step 4: Calculate ROC curve;
[…]
12:Step 5: Calculate AUROC;
[…]
14:Step 6: Calculate Precision-Recall curve;
[…]
16:Step 7: Calculate AUPRC;
[…]
AUROC (Area Under the Receiver Operating Characteristic Curve) and AUPRC (Area Under the Precision-Recall Curve) are evaluation metrics commonly used to assess the performance of binary classification models, providing insights into their ability to distinguish between positive and negative instances. In the context of evaluating uncertainty estimation in machine learning models, these metrics quantify how well the model separates iD samples from OoD samples. AUROC is derived from the Receiver Operating Characteristic (ROC) curve, which illustrates the trade-off between True Positive Rate (TPR) and False Positive Rate (FPR) across various classification thresholds.
The AUROC value represents the area under this curve and ranges from 0 to 1, with higher values indicating better discriminative performance. AUPRC is based on the Precision-Recall curve, which plots Precision against Recall at different classification thresholds. Precision measures the accuracy of positive predictions, while Recall quantifies the ability to capture all positive instances. AUPRC calculates the area under this curve and provides a complementary perspective, particularly valuable when dealing with imbalanced datasets.
@@
Figs. 13 and 13 plot each model’s performance, illustrating the trade-offs between True Positive Rate and False Positive Rate (ROC curve), Precision and Recall (PRC curve) for CIFAR-10 (Fig. 13) and MNIST (Fig. 13). The left two plots depict the AUROC curves, where the blue curve representing RS-NN outperforms others, indicating superior discrimination between true positive and false positive rates. Similarly, on the two right plots displaying Precision-Recall Curve (PRC), RS-NN exhibits the highest curve, emphasizing its precision and recall performance. These results showcase RS-NN’s effectiveness in distinguishing in-distribution and out-of-distribution samples.
In real-world scenarios, encountering OoD instances is inevitable, making reliable OoD detection essential for safety-critical applications to avoid erroneous decisions on unfamiliar data. Recognising unfamiliar data signals the model about situations beyond its training, allowing it to acknowledge its own limitations and ignorance, which in turn enhances its uncertainty estimation.
Appendix D Implementation Details
Laplace Bridge Bayesian approximation (LB-BNN) (Hobbhahn et al., 2022) uses the Laplace Bridge to map efficiently between Gaussian and Dirichlet distributions and enhance computational efficiency. In our experiments, we obtain multiple samples from the LB-BNN posterior and average these samples using Bayesian Model Averaging (BMA). Function-Space Variational Inference (FSVI) (Rudner et al., 2022) utilizes variational inference to derive an approximation of the posterior distribution over the function space. FSVI proposes approximating distributions over functions as Gaussian by linearizing their mean parameters and derived a tractable and well-defined variational posterior. Deep Ensembles (DEs) (Lakshminarayanan et al., 2017) and Epistemic Neural Networks (ENNs) (Osband et al., 2024) are ensemble methods that train ensembles of models to efficiently estimate uncertainty. The mean and variance of ensemble predictions are calculated and the mean is softmaxed to obtain predictions from DE and ENN. The training hyperparameters for all models are outlined in Tab. 5.
Table 5: Training Hyperparameters for All Models
Parameter Value
Architecture ResNet50 (excluding top classification layer)
Additional Dense Layers 1024 and 512 neurons (ReLU activation)
Output Layer Activation (RS-NN) Sigmoid (multi-label classification)
Output Layer Activation (Other Models) Softmax (multi-class classification)
Learning Rate Initial: 1e-3
Learning Rate Scheduler Decrease by 0.1 at epochs 80, 120, 160, 180
Training Epochs 200
Batch Size 128
Optimizer RS-NN: Adam, LB-BNN: Adam, ENN: Adam, DE: Adam, FSVI: SGD
Training Dataset Sizes
CIFAR-10: 40,000
CIFAR-100: 40,000
MNIST: 50,000
Intel Image: 13,934
ImageNet: 1,172,498
Testig Dataset Sizes
CIFAR-10: 10,000
CIFAR-100: 10,000
MNIST: 10,000
Intel Image: 3,000
ImageNet: 2,000
Testing Samples for OoD Datasets 10,000 (except Intel Image: 3,000)
Input Image Size 224 × 224
Data Augmentation Random horizontal/vertical shifts (magnitude 0.1), horizontal flips
GPU NVIDIA A100 80GB
We report the training times (in minutes) for all models on the CIFAR-10 dataset in Tab. 6. Tab. 6 shows that FSVI has the longest training time, followed by RNN and DE. Compared to LB-BNN, RS-NN has an additional 5 minutes to the training duration, while CNN has the shortest training time. However, CNNs do not provide uncertainty estimation, and RS-NN has more reliable uncertainty and OoD metrics in comparison to LB-BNN.
The significant training time required for Deep Ensembles (DE) often does not justify their uncertainty estimates (Abe et al., 2022; Rahaman et al., 2021). Despite the computational cost, DE is sometimes only as good as other uncertainty estimation models, which can be trained much faster. Other approaches, like methods with calibration techniques (e.g., temperature scaling or Bayesian approaches), can provide similar performance without the substantial overhead that DE entails.